At what pH will phenobarbital begin to separate from its solution? Phenobarbital is a weak acid.
Initial conc. 0.0393M
Ionized conc. 0.0043M
Ka 3.9x10^-8

ROUND OFF TO 2 DECIMAL PLACES.
1
Expert's answer
2020-10-29T05:23:43-0400

Solution.

"3.9\times 10^{-8} = \frac{x^2}{(0.0393-x)}"

"x = 3.91 \times 10^{-5}"

"pH = -lg[H^+] = 4.41"

Answer:

4.41